# Constructing Concise Characteristic Samples for Acceptors of Omega Regular Languages

### v3

@@ 16.4. Testing membership in $\mathbb{IMA}$
A DMA $\mathcal{U}=\langle\Sigma,Q,q_{\iota},\delta,{\mathcal{F}}\rangle$
If ${\llbracket}\mathcal{U}{\rrbracket}\in\mathbb{IMA}$ then return an IMA accepting ${\llbracket}\mathcal{U}{\rrbracket}$, else return “no”
@@ A DMA $\mathcal{U}=\langle\Sigma,Q,q_{\iota},\delta,{\mathcal{F}}\rangle$
The algorithm calls the $\mathbf{RightCon}$ algorithm and also the DMA equivalence algorithm from Thm. 14.6, which run in polynomial time. When there is a witness $u(v)^{\omega}$ accepted by $\mathcal{U}$, there is a set $F\in{\mathcal{F}}$ whose image in $\mathcal{M}$ is added to ${\mathcal{F}}^{\prime}$, so there can be no more such witnesses than the number of sets in ${\mathcal{F}}$. After this, there must be a successful equivalence test or a witness rejected by $\mathcal{U}$, either of which terminates the while loop. Thus, the overall running time is polynomial in the size of $\mathcal{U}$.
@@ 16.6. Efficient teachability of the informative classes
Let $\mathcal{A}$ be an acceptor of type DXA for $X\in\{B,C,P,R,S,M\}$. By Theorems 16.1, 16.2, 16.6, 16.7, there are polynomial time algorithms that return an equivalent acceptor $\mathcal{A^{\prime}}$ in IXA if such exists and “no” otherwise. By Thm. 10.6, given an acceptor $\mathcal{A}^{\prime}$ of type IBA, ICA, IPA, IRA, ISA, or IMA, the characteristic sample $T_{L}$ for $\mathcal{A}$ may be computed in polynomial time in the size of $\mathcal{A}$. It follows that the classes $\mathbb{IMA}$, $\mathbb{IBA}$, $\mathbb{ICA}$, $\mathbb{IPA}$, $\mathbb{IRA}$ and $\mathbb{ISA}$ are efficiently teachable. ∎
@@ 17. Discussion
Focusing on the classes of informative languages, $\mathbb{IBA}$, $\mathbb{ICA}$, $\mathbb{IPA}$, $\mathbb{IRA}$, $\mathbb{ISA}$, and $\mathbb{IMA}$, we have shown that they are efficiently teachable. To obtain these results we have given new polynomial time algorithms to test inclusion and equivalence for DBAs, DCAs, DPAs, DRAs, DSAs, and DMAs. We have given a polynomial time algorithm to compute the right congruence automaton $\mathcal{M}_{\sim_{L}}$ for a language $L$ specified by a DBA, DCA, DPA, DRA, DSA, or DMA. This yields a polynomial time algorithm to test whether an acceptor $\mathcal{A}$ of type DBA is of type IBA, and similarly for acceptors of types DCA, DPA, DRA, DSA, and DMA. Moreover, we have given a polynomial time algorithm to test whether an acceptor $\mathcal{A}$ of type DBA accepts a language in the class $\mathbb{IBA}$, and similarly for acceptors of types DCA, DPA, DRA, DSA and DMA.
The questions of whether the full deterministic classes $\mathbb{DBA}$, $\mathbb{DCA}$, $\mathbb{DPA}$, $\mathbb{DRA}$, $\mathbb{DSA}$, and $\mathbb{DMA}$ are efficiently teachable or identifiable in the limit using polynomial time and data remain open. We note that Bohn and Löding [BL22, BL23] have recently obtained interesting results for passive learning of $\mathbb{DBA}$ and $\mathbb{DPA}$, but from characteristic samples of cardinality that may be exponential in the size of the minimal representations in the worst case. Thus these results do not settle the question of whether these classes are efficiently teachable or identifiable in the limit using polynomial time and data. Another intriguing open question is whether the classes $\mathbb{IBA}$, $\mathbb{ICA}$, $\mathbb{IPA}$, $\mathbb{IRA}$, $\mathbb{ISA}$, and $\mathbb{IMA}$ can be learned by polynomial time algorithms using membership and equivalence queries. However, as shown by Bohn and Löding [BL21], this question is not easier than whether the corresponding deterministic classes can be learned by polynomial time algorithms using membership and equivalence queries.

### v4

@@ 16.4. Testing membership in $\mathbb{IMA}$
A DMA $\mathcal{U}=\langle\Sigma,Q,q_{\iota},\delta,{\mathcal{F}}\rangle$.
If ${\llbracket}\mathcal{U}{\rrbracket}\in\mathbb{IMA}$ then return an IMA accepting ${\llbracket}\mathcal{U}{\rrbracket}$, else return “no”.
@@
The algorithm calls the $\mathbf{RightCon}$ algorithm and also the DMA equivalence algorithm from Theorem 14.6, which run in polynomial time. When there is a witness $u(v)^{\omega}$ accepted by $\mathcal{U}$, there is a set $F\in{\mathcal{F}}$ whose image in $\mathcal{M}$ is added to ${\mathcal{F}}^{\prime}$, so there can be no more such witnesses than the number of sets in ${\mathcal{F}}$. After this, there must be a successful equivalence test or a witness rejected by $\mathcal{U}$, either of which terminates the while loop. Thus, the overall running time is polynomial in the size of $\mathcal{U}$.
@@ 16.6. Efficient teachability of the informative classes
Let $\mathcal{A}$ be an acceptor of type DXA for $X\in\{B,C,P,R,S,M\}$. By Theorems 16.1, 16.2, 16.6, 16.7, there are polynomial time algorithms that return an equivalent acceptor $\mathcal{A^{\prime}}$ in IXA if such exists and “no” otherwise. By Theorem 10.6, given an acceptor $\mathcal{A}^{\prime}$ of type IBA, ICA, IPA, IRA, ISA, or IMA, the characteristic sample $T_{L}$ for $\mathcal{A}$ may be computed in polynomial time in the size of $\mathcal{A}$. It follows that the classes $\mathbb{IMA}$, $\mathbb{IBA}$, $\mathbb{ICA}$, $\mathbb{IPA}$, $\mathbb{IRA}$ and $\mathbb{ISA}$ are efficiently teachable. ∎
@@ 17. Discussion
Focusing on the classes of informative languages, $\mathbb{IBA}$, $\mathbb{ICA}$, $\mathbb{IPA}$, $\mathbb{IRA}$, $\mathbb{ISA}$ and $\mathbb{IMA}$, we have shown that they are efficiently teachable. To obtain these results we have given new polynomial time algorithms to test inclusion and equivalence for DBAs, DCAs, DPAs, DRAs, DSAs and DMAs. We have given a polynomial time algorithm to compute the right congruence automaton $\mathcal{M}_{\sim_{L}}$ for a language $L$ specified by a DBA, DCA, DPA, DRA, DSA, or DMA. This yields a polynomial time algorithm to test whether an acceptor $\mathcal{A}$ of type DBA is of type IBA, and similarly for acceptors of types DCA, DPA, DRA, DSA and DMA. Moreover, we have given a polynomial time algorithm to test whether an acceptor $\mathcal{A}$ of type DBA accepts a language in the class $\mathbb{IBA}$, and similarly for acceptors of types DCA, DPA, DRA, DSA and DMA.
The questions of whether the full deterministic classes $\mathbb{DBA}$, $\mathbb{DCA}$, $\mathbb{DPA}$, $\mathbb{DRA}$, $\mathbb{DSA}$ and $\mathbb{DMA}$ are efficiently teachable or identifiable in the limit using polynomial time and data remain open. We note that Bohn and Löding [BL22, BL23] have recently obtained interesting results for passive learning of $\mathbb{DBA}$ and $\mathbb{DPA}$, but from characteristic samples of cardinality that may be exponential in the size of the minimal representations in the worst case. Thus these results do not settle the question of whether these classes are efficiently teachable or identifiable in the limit using polynomial time and data. Another intriguing open question is whether the classes $\mathbb{IBA}$, $\mathbb{ICA}$, $\mathbb{IPA}$, $\mathbb{IRA}$, $\mathbb{ISA}$ and $\mathbb{IMA}$ can be learned by polynomial time algorithms using membership and equivalence queries. However, as shown by Bohn and Löding [BL21], this question is not easier than whether the corresponding deterministic classes can be learned by polynomial time algorithms using membership and equivalence queries.